OpenStudy (anonymous):

solve the system of equations x+2y=6 x^2+y^2=8

OpenStudy (anonymous):

I got 24 plus or minus radical 1136 all over 10 and I feel like what a I suppose to do with that.

OpenStudy (anonymous):

x = 6 - 2y (6 - 2y)^2 + y^2 = 8 4y^2 -24y + 36 + y ^2 = 8 5y^2 -24y + 36 = 8 5y^2 - 24y + 28 = 0 y = 2 y = 14/5 sub those back in to the linear

OpenStudy (anonymous):

I did everything u did but from the last line I went and plugged in the quadratic equation where did u get y=2 and y =14/5?

OpenStudy (anonymous):

I still need help, im not understanding this.

OpenStudy (anonymous):

solving 5y^2 - 24y + 28 = 0 you get y = 2, 14/5

OpenStudy (anonymous):

ya but what do u mean solving that, what did u do to solve it because I tried the quadratic and got this weird number.

OpenStudy (anonymous):

solve for the zeroes, what else?

OpenStudy (anonymous):

use quad formula, factor, anything you want.

OpenStudy (anonymous):

ugh forget it.